OpenStudy (anonymous):

What is the slope of the line that passes through the given points? (2, 12) and (6, 11)

hero (hero):

1. Write the equation of the line: y = mx + b 2. Insert the points unto the line: 12 = 2m + b 11 = 6m + b 3. Isolate b in each: b = 12 - 2m b = 11 - 6m 4. Set b = b 12 - 2m = 11 - 6m 5. Solve for m: 6m - 2m = 11 - 12 4m = -1 m = -1/4 The slope of the line passing through both points is -1/4

OpenStudy (anonymous):

Or you can use slope formula: \(\large m = \dfrac{y_2 - y_1}{x_2 - x_1}\)

hero (hero):

You could, but, that would require memorizing another formula. I'm showing here that you only need to know y = mx + b to figure out what the slope is.

OpenStudy (anonymous):

Memorizing formulas make math much easier. Just saying.

hero (hero):

Let's say, hypothetically, you are taking an exam, and suddenly forget the slope formula. Well, if you forget the slope formula, but still remember the equation of the line, you can still get the correct answer. Just saying.

hero (hero):

I could type up the solution faster than you could write up your slope formula.

OpenStudy (anonymous):

Good for you. Let's not argue and let OP decides what to do.

hero (hero):

I'm pretty sure she already decided.

OpenStudy (anonymous):

thanks @Hero you are my hero hehe i love you